OpenStudy (shrutipande9):

find stationary points and examine for max & min. f(x,y)=x^2+y^2+3xy

OpenStudy (chmvijay):

@ganeshie8 @hartnn @wio @Directrix @nubeer can u help her

OpenStudy (john_es):

You need to find the gradient first, \[\nabla f(x,y)=(2x+3y,2y+3x)\]Then, find the candidate points, \[\nabla f(x,y)=(2x+3y,2y+3x)=(0,0)\Rightarrow(x,y)=(0,0)\]Now, you must find which type of point you have. You need the Hessian (may be you can use another criteria), \[H(x,y)=\left(\begin{matrix} \frac{\partial^2 f}{\partial x^2} & \frac{\partial^2 f}{\partial x\partial y} \\ \frac{\partial^2 f}{\partial y\partial x} & \frac{\partial^2 f}{\partial y^2}\end{matrix}\right)=\left(\begin{matrix} 2 & 3 \\ 3& 2\end{matrix}\right)\] Use the Sylvester cryteria, and you will find that you have a saddle point in (0,0) but no maxima or minima.

OpenStudy (shrutipande9):

okay thank u sooo mch!

OpenStudy (john_es):

;)